# Tree Ensembles for Contextual Bandits

### v2

@@ Abstract
We propose a novel framework for contextual multi-armed bandits based on tree ensembles. Our framework integrates two widely used bandit methods, Upper Confidence Bound and Thompson Sampling, for both standard and combinatorial settings. We demonstrate the effectiveness of our framework via several experimental studies, employing both XGBoost and random forest, two popular tree ensemble methods. Compared to state-of-the-art methods based on decision trees and neural networks, our methods exhibit superior performance in terms of both regret minimization and computational runtime, when applied to benchmark datasets and the real-world application of navigation over road networks.
@@ 1 Introduction
Stochastic multi-armed bandits (MABs) (Slivkins, 2019) provide a principled framework for making optimal sequences of decisions under uncertainty. As discussed by Zhu and Van Roy (2023), MABs constitute a vital component of modern recommendation systems for efficient exploration, among many other applications. An important variant known as the contextual multi-armed bandit incorporates additional contextual / side information into the decision-making process, allowing for more personalized or adaptive action selection. Following the recent success of (deep) neural networks in solving various machine learning tasks, several methods building on such models have been suggested for finding functional relationships between the context and outcomes of actions to aid the decision-making process (Zhou et al., 2020; Zhang et al., 2021; Zhu and Van Roy, 2023; Osband et al., 2021). Although delicate, these methods can be impractical, time-consuming, and computationally expensive.
In light of these challenges, we propose bandit algorithms that utilize tree ensemble (TE) methods (Hastie et al., 2009), like gradient-boosted decision trees (GBDT) and random forests, to comprehend the contextual features, combined with the most popular bandit methods Upper Confidence Bound (UCB) and Thompson Sampling (TS) as exploration strategies. Compared with both other tree-based and neural network methods, our methods, called TEUCB and TETS, yield superior results on UCI benchmark datasets. Furthermore, our methods benefit from more effective learning with less computational overhead. We furthermore extend the framework to the combinatorial contextual setting, a variant of bandits that deals with complex combinatorial sets of arms, and investigate them on the important real-world application of efficient navigation in road networks.
@@ 1.1 Related Work
The concept of applying decision trees to contextual bandit problems has been studied to a limited extent. Elmachtoub et al. (2017) suggest using a single decision tree for each arm available to the agent. Their method is referred to as TreeBootstrap, as the concept of bootstrapping is employed in order to resemble the Thompson Sampling process. One issue with their method is that it might have trouble reaching the accuracy and robustness of tree ensemble methods, especially when estimating complex reward functions. The possibility of applying random forests in their framework is discussed, but no concrete methods or experimental results are presented. Furthermore, storing and training one model per arm does not scale, especially not with large action spaces (e.g., in the combinatorial setting), and could potentially lead to excessive exploration for dynamic arm sets, as it cannot attend to what it has learned from the contexts of other arms.
@@
In contrast, Féraud et al. (2016) employ a tree ensemble but only uses decision trees of unit depth also known as decision stumps. While tree ensembles with a large number of decision stumps tend to perform well regarding accuracy, limited variance, and increased robustness, restricting the depth in such an extreme way may not be adequate when addressing complex tasks (Hastie et al., 2009). The experimental studies by Hastie et al. (2009) indicate that tree depths of 4 to 8 work well for boosting. Additionally, Féraud et al. (2016) only investigate the bandit algorithm known as successive elimination, which usually leads to less efficient exploration compared to UCB and TS, and thus is used less commonly. Moreover, the method, called Bandit Forest, requires binary encoded contextual features. Comparing these two tree-based bandit methods, Elmachtoub et al. (2017) experimentally show that TreeBootstrap tends to outperform Bandit Forest in practice.
@@
It is worth noting that neither Elmachtoub et al. (2017) nor Féraud et al. (2016) consider combinatorial contextual bandits. Apart from these two methods, the problem has primarily been addressed using other machine learning models. When the reward function is linear w.r.t. the context, LinUCB (Li et al., 2010) and Linear Thompson Sampling (LinTS) (Agrawal and Goyal, 2013) have demonstrated good performance. In more complicated cases, the linear assumption no longer holds, and thus more expressive models are needed.
@@
In recent years, neural contextual bandits have attracted a lot of attention. As the name suggests, these methods use (deep) neural networks to model the expected rewards for some set of observed contextual features. Out of several proposed neural contextual bandit algorithms, NeuralUCB (Zhou et al., 2020) and NeuralTS (Zhang et al., 2021), in particular, have been shown to perform well, while also providing theoretical regret bounds. On the negative side, due to how these methods estimate the uncertainty in predictions, they rely on inverting a matrix of size equal to the number of network parameters, which is usually computationally expensive and time-consuming, especially when the prediction tasks require large neural networks.
@@
Zhu and Van Roy (2023) provide a thorough review of different neural bandit methods and suggests a method based on epistemic neural networks (Osband et al., 2021) for more efficient uncertainty modeling, showing its advantages in terms of regret minimization and computational efficiency. Despite their promising results, we argue that neural networks may not necessarily be the most appropriate models for contextual bandits. For instance, even with epistemic neural networks, training is still resource-intensive. On the other hand, there is a large body of work demonstrating the effectiveness of tree ensemble methods for both regression and classification tasks in the standard supervised learning paradigm (Borisov et al., 2022; Gorishniy et al., 2021; Grinsztajn et al., 2022), and thereby, we believe that their extension to contextual bandits can be of great potential.
@@ 2.1 Multi-Armed Bandit Problem
The multi-armed bandit (MAB) problem is a sequential decision-making problem under uncertainty, where the decision maker, i.e., the agent, interactively learns from the environment over a horizon of $T$ time steps (interactions). At each time step $t\leq T$, the agent is presented with a set of $K$ actions $\mathcal{A}$, commonly referred to as arms. Each action $a\in\mathcal{A}$ has an (initially) unknown reward distribution with expected value $\mu_{a}$. The objective is to maximize the cumulative reward over the time horizon $T$, or more commonly (and equivalently) to minimize the cumulative regret, which is calculated as
$R(T)\triangleq\sum^{T}_{t=1}(\mu^{*}-\mu_{a_{t}}),$ (1)
where $\mu^{*}$ is the expected reward of the optimal arm, and $a_{t}$ is the arm selected at time step $t$. It is important to note that the agent only receives feedback from the selected arm / action. More specifically, when selecting $a_{t}$ the agent receives a reward $r_{t,a_{t}}$ which is sampled from the underlying reward distribution. As previously mentioned, the agent’s objective is to minimize the cumulative regret, not to learn the complete reward function for each arm. Thus, the agent has to balance delicately between exploration and exploitation.
@@ 2.2 Contextual Bandit Problem
The contextual multi-armed bandit problem is an extension of the classical MAB problem described in Section 2.1. The agent is, at each time step $t$, presented with a context vector $\mathbf{x}_{t,a}\in\mathbb{R}^{d}$ for each action $a\in\mathcal{A}$. For example, a recommendation system could encode a combination of user-related and action-specific features into the context vector $\mathbf{x}_{a}$. Then, the expected reward of action $a$ at time $t$ is given by an (unknown) function of the context $f:\mathbb{R}^{d}\rightarrow\mathbb{R}$, such that $\mathbb{E}[r_{t,a}]=f(\mathbf{x}_{t,a})$. Learning to generalize the relationship between the contextual features and the expected reward is crucial for effective learning and minimizing the cumulative regret.
@@ 2.3 Combinatorial Bandits
Combinatorial multi-armed bandits (CMAB) (Cesa-Bianchi and Lugosi, 2012) deal with problems where at each time step $t$, a subset $\mathcal{S}_{t}$ (called a super arm) of the set of all base arms $\mathcal{A}$ is selected, instead of an individual arm. The reward of a super arm $\mathcal{S}$ depends on its constituent base arms. When the reward for a super arm is given as a function (e.g., sum) of feedback from its individual base arms (observable if and only if the super arm is selected), it is referred to as semi-bandit feedback. As previously, the evaluation metric is the cumulative regret, and for the combinatorial semi-bandit setting (with the sum of base arm feedback as super arm reward) it can be calculated as
@@ 3 Proposed Algorithms
Despite the potential benefits, these types of models have not been studied much for contextual bandits. To the best of our knowledge, there are no previously known works that in a principled way combine tree ensemble models with UCB, which is one of the most effective strategies known for handling the exploration-exploitation dilemma. Further, works that combine tree ensemble methods with Thompson Sampling are limited. We address this research gap and introduce a novel approach for the contextual MAB problem that combines any tree ensemble method with a natural adaption of these two prominent exploration schemes.
Within our framework, we empirically investigate the XGBoost and random forest algorithms and demonstrate their auspicious performance on several benchmark tasks, as well as a combinatorial contextual MAB problem for efficient navigation over a stochastic road network in Section 4. However, we emphasize that our bandit algorithms are sufficiently general to be employed with any decision tree ensemble.
3.1 Weak Learners
@@
In our notation, a tree ensemble regressor $F\left(\cdot\>;\{(\cdot,\cdot)\}\right)$ is a collection of decision trees, where the trees are collectively fitted to a set of training data points with corresponding target values $\{(\mathbf{x},r)\}$. The fitting procedure can differ between different types of tree ensemble methods, which are often divided into two different categories depending on which fitting procedure is employed, i.e., bagging or boosting. In both variants, the prediction of each tree is determined by assigning the training data points to distinct leaves, where all data points in the same leaf resemble each other to some degree, based on the features attended to in the splitting criteria associated with the ancestor nodes of the leaf. Hence, when fitted to the data, each tree accepts an input vector $\mathbf{x}$ which it assigns to one of its leaves.
Every leaf is associated with three values that depend on the samples from the training data assigned to it. We denote the output of each tree as $o$, which is its contribution to the total ensemble prediction. The number of training data points assigned to each leaf is denoted $c$. Finally, the sample variance in the output proposed by each individual training data point assigned to the leaf is denoted $s^{2}$. Note that, for the sample variance to exist, we must have at least two samples. This is ensured by requiring the tree to be built such that no leaf has less than two training data points assigned to it. Then, the tree ensemble constructs its total prediction $p$ by summing up all of the $N$ individual trees’ outputs, such that
$p=\sum_{i=1}^{N}o_{i},$ (3)
where the outputs $o_{i}$, in turn, are averages of the output suggested by each training sample based on its target value. We use this definition of $p$ to unify the notation of all kinds of tree ensembles. Hence, for averaging models, such as random forest, we assume that the leaf values $o_{i}$ are already weighted by $1/N$.
@@ 3.2 Uncertainty Modeling
Firstly, we assume that the output $o_{n}$ of the $n$’th decision tree, given a context $\mathbf{x}$, is an arithmetic average of $c_{n}$ independent and identically distributed random variables with finite mean $\mu_{n}$ and variance $\sigma_{n}^{2}$. By this assumption, $o_{n}$ is itself a random variable with mean $\mu_{n}$ and variance $\frac{\sigma_{n}^{2}}{c_{n}}$. Also, we can approximate $\mu_{n}$ and $\sigma_{n}^{2}$ by the sample mean and variance respectively, based on the training data points assigned to the leaf. Moreover, the central limit theorem (CLT) (see e.g., Dodge, 2008) ensures that, as $c_{n}\to\infty$, the distribution of $o_{n}$ tends to a Gaussian distribution.
Secondly, if we also assume the output of each one of the $N$ trees in the ensemble to be independent of every other tree, we have that the total prediction $p$, which by Equation 3 is a sum of $N$ random variables, is normally distributed with mean $\mu=\sum_{i=1}^{N}\mu_{i}$ and variance $\sigma^{2}=\sum_{i=1}^{N}\sigma_{i}^{2}$.
@@ 3.3 Tree Ensemble Upper Confidence Bound
UCB methods act under the principle of optimism in the face of uncertainty, and have established themselves as one of the most prominent approaches to handling exploration in bandit problems. A classic example is the UCB1 algorithm (Auer et al., 2002) which builds confidence bounds around the expected reward from each arm depending on the fraction of times they have been played, and for which there are proven upper bounds on the expected regret. One disadvantage of the method, however, is that it does not take the variance of the observed rewards into account, which may lead to a sub-optimal exploration strategy in practice. In light of this, Auer et al. (2002) further proposed the UCB1-Tuned and UCB1-Normal algorithms, which extend UCB1 by including variance estimates, and demonstrate better performance experimentally. The main difference between the two extended versions is that UCB1-Tuned assumes Bernoulli-distributed rewards, while UCB1-Normal is constructed for Gaussian rewards. At each time step $t$, UCB1-Normal selects the arm $a$ with the maximal upper confidence bound $U_{t,a}$, calculated as
$U_{t,a}\leftarrow\tilde{\mu}_{t,a}+\sqrt{16\tilde{\sigma}^{2}_{t,a}\frac{\ln(t-1)}{n_{t,a}}},$ (4)
where $n_{t,a}$ is the number of times arm $a$ has been played so far, and $\tilde{\mu}_{t,a}$ and $\tilde{\sigma}^{2}_{t,a}$ are the sample mean and variance of the corresponding observed rewards, respectively. For the sample variance to be defined for all arms, they must have been played at least twice first.
By the assumptions we have made on the tree ensembles, the predictions of the expected rewards will be approximately Gaussian. Therefore, we propose an algorithm called Tree Ensemble Upper Confidence Bound (TEUCB) in Algorithm 1, which draws inspiration from UCB1-Normal, and is suitable for both Gaussian and Bernoulli bandits. As seen in lines 22 and 23 of Algorithm 1, the selection rule of TEUCB closely resembles that of UCB1-Normal, but is constructed specifically for contextual MABs where the arms available in each time step are characterized by their context vectors. Therefore, TEUCB considers each individual contribution from all data points in a leaf as a sample of that leaf’s output distribution, which is demonstrated in line 13. The total prediction for a context $\mathbf{x}_{t,a}$ is subsequently computed as the sum of sampled contributions from each leaf that $\mathbf{x}_{t,a}$ is assigned to in the ensemble (lines 15, 16, 17).
Algorithm 1 Tree Ensemble Upper Confidence Bound
1:  Input: Number of rounds $T$, number of initial random selection rounds $T_{I}$, number of trees in ensemble $N$, exploration factor $\nu$, tree ensemble regressor $F(\cdot\>;\{\cdot,\cdot\})$.
@@
7:     Observe contexts $\{\mathbf{x}_{t,a}\}_{a=1}^{K}$
8:     for $a=1$ to $K$ do
9:        Initialize arm parameters:$\tilde{\mu}_{t,a}\leftarrow 0,\quad\tilde{\sigma}^{2}_{t,a}\leftarrow 0,\quad c_{t,a}\leftarrow 0$
10:        for n = 1 to $N$ do
11:           Assign leaf values:
12:           $(o_{t,a,n},s_{t,a,n},c_{t,a,n})$ $\leftarrow F_{n}\left(\mathbf{x}_{t,a};\left\{(\mathbf{x}_{i,a_{i}},r_{i,a_{i}})\right\}_{i=1}^{t-1}\right)$
13:           Estimate expected partial reward distribution: $\tilde{\mu}_{t,a,n}\leftarrow o_{t,a,n},\quad\tilde{\sigma}^{2}_{t,a,n}\leftarrow\frac{s^{2}_{t,a,n}}{c_{t,a,n}}$
14:           Increment arm parameters:
15:           $\tilde{\mu}_{t,a}\leftarrow\tilde{\mu}_{t,a}+\tilde{\mu}_{t,a,n}$
16:           $\tilde{\sigma}^{2}_{t,a}\leftarrow\tilde{\sigma}^{2}_{t,a}+\tilde{\sigma}^{2}_{t,a,n}$
17:           $c_{t,a}\leftarrow c_{t,a}+c_{t,a,n}$
18:        end for
19:        Calculate UCB:
20:        $U_{t,a}\leftarrow\tilde{\mu}_{t,a}+\sqrt{\nu^{2}\tilde{\sigma}^{2}_{t,a}\frac{\ln(t-1)}{c_{t,a}}}$
21:     end for
22:     $a_{t}\leftarrow\text{argmax}_{a}U_{t,a}$
23:     Play $a_{t}$ and observe reward $r_{t,a_{t}}$
24:  end for

### v3

@@ Abstract
We propose a new framework for contextual multi-armed bandits based on tree ensembles. Our framework adapts two widely used bandit methods, Upper Confidence Bound and Thompson Sampling, for both standard and combinatorial settings. As part of this framework, we propose a novel method of estimating the uncertainty in tree ensemble predictions. We further demonstrate the effectiveness of our framework via several experimental studies, employing XGBoost and random forests, two popular tree ensemble methods. Compared to state-of-the-art methods based on decision trees and neural networks, our methods exhibit superior performance in terms of both regret minimization and computational runtime, when applied to benchmark datasets and the real-world application of navigation over road networks.
@@ 1 Introduction
Stochastic multi-armed bandits (MABs) (see Slivkins, 2019) provide a principled framework for making optimal sequences of decisions under uncertainty. As discussed by Zhu and Van Roy (2023), MABs constitute a vital component of modern recommendation systems for efficient exploration, among many other applications. An important variant known as the contextual multi-armed bandit incorporates additional contextual / side information into the decision-making process, allowing for more personalized or adaptive action selection. Following the recent success of (deep) neural networks in solving various machine learning tasks, several methods building on such models have been suggested for finding functional relationships between the context and outcomes of actions to aid the decision-making process (Zhou et al., 2020; Zhang et al., 2021; Zhu and Van Roy, 2023; Osband et al., 2021; Hoseini et al., 2022). Even though sophisticated, these methods can be impractical, time-consuming, and computationally expensive.
In light of these challenges, we propose bandit algorithms that utilize tree ensemble (TE) methods (Hastie et al., 2009), like gradient-boosted decision trees (GBDT) and random forests, to comprehend the contextual features, combined with the most popular bandit methods Upper Confidence Bound (UCB) and Thompson Sampling (TS) as exploration strategies. Compared with other tree-based and neural network methods, our methods, called TEUCB and TETS, yield superior results on UCI benchmark datasets. Additionally, our methods benefit from more effective learning with less computational overhead on most problem instances, compared to existing methods that use similarly expressive machine learning models. We furthermore extend the framework to the combinatorial contextual setting, a type of bandit that deals with complex combinatorial sets of arms, and investigate it on the important real-world application of efficient navigation in road networks. In this work, we focus on the practical applicability of the proposed framework in different settings and do not provide a theoretical analysis of the regret.
@@ 1.1 Related Work
The concept of applying decision trees to contextual bandit problems has been studied to a limited extent. Elmachtoub et al. (2017) suggest using a single decision tree for each arm available to the agent. Their method is referred to as TreeBootstrap, as the concept of bootstrapping is employed in order to resemble the Thompson Sampling process. One issue with their method is its limited ability to achieve the accuracy and robustness of tree ensemble methods, particularly when estimating complex reward functions. The possibility of applying random forests in their framework is discussed, but no concrete methods or experimental results are presented. Furthermore, storing and training one tree model per arm does not scale well, especially not with large action spaces (e.g., in the combinatorial setting), and could potentially lead to excessive exploration for dynamic arm sets, as it cannot attend to what it has learned from the contexts of other arms. In contrast, Féraud et al. (2016) employ a tree ensemble but only use decision trees of unit depth, also known as decision stumps. While tree ensembles with a large number of decision stumps tend to perform well regarding accuracy, lower variance, and increased robustness, restricting the depth in such an extreme way may not be adequate when addressing complex tasks (Hastie et al., 2009). The experimental studies by Hastie et al. (2009) indicate that tree depths of 4 to 8 work well for boosting. Additionally, Féraud et al. (2016) only investigate the bandit algorithm known as successive elimination, which usually leads to less efficient exploration compared to UCB and TS, and thus is used less commonly. Moreover, the method, called Bandit Forest, requires binary encoded contextual features. Comparing these two tree-based bandit methods, Elmachtoub et al. (2017) experimentally show that TreeBootstrap tends to outperform Bandit Forest in practice. It should be noted that neither Elmachtoub et al. (2017) nor Féraud et al. (2016) consider combinatorial contextual bandits.
@@
Apart from the two methods mentioned above, the problem has primarily been addressed using other machine learning models. When the reward function is linear w.r.t. the context, LinUCB (Li et al., 2010; Abbasi-Yadkori et al., 2011) and Linear Thompson Sampling (LinTS) (Agrawal and Goyal, 2013) have demonstrated good performance. In more complicated cases, the linear assumption no longer holds, and thus more expressive models are needed.
@@
In recent years, neural contextual bandits have attracted a lot of attention. As the name suggests, these methods use (deep) neural networks to model the expected rewards for some observed contextual features. Out of several proposed neural contextual bandit algorithms, NeuralUCB (Zhou et al., 2020) and NeuralTS (Zhang et al., 2021), in particular, have been shown to perform well, while also providing theoretical regret bounds. On the negative side, due to how these methods estimate the uncertainty in predictions, they rely on inverting a matrix of size equal to the number of network parameters, which is usually computationally expensive and time-consuming, especially when the prediction tasks require large neural networks.
@@
Zhu and Van Roy (2023) provide a thorough review of different neural bandit methods and suggest a method based on epistemic neural networks (Osband et al., 2021) for more efficient uncertainty modeling, showing its advantages in terms of regret minimization and computational efficiency. Despite their promising results, we argue that neural networks may not necessarily be the most appropriate models for contextual bandits. For instance, even with epistemic neural networks, training is still resource-intensive. On the other hand, there is a large body of work that shows the effectiveness of tree ensemble methods for regression and classification tasks in the standard supervised learning setting (Borisov et al., 2022; Gorishniy et al., 2021; Grinsztajn et al., 2022), and therefore we believe that their extension to contextual bandits can have great potential.
@@ 2.1 Multi-Armed Bandit Problem
The multi-armed bandit (MAB) problem is a sequential decision-making problem under uncertainty, where the decision maker, i.e., the agent, interactively learns from the environment over a horizon of $T$ time steps (interactions). At each time step $t\leq T$, the agent is presented with a set of $K$ actions $\mathcal{A}$, commonly referred to as arms. Each action $a\in\mathcal{A}$ has an (initially) unknown reward distribution with expected value $\mu_{a}$. The objective is to maximize the cumulative reward over the time horizon $T$, or more commonly (and equivalently) to minimize the cumulative regret, which is defined as
$R(T)\triangleq\sum^{T}_{t=1}(\mu_{a^{*}}-\mu_{a_{t}}),$ (1)
where $\mu_{a^{*}}$ is the expected reward of the optimal arm $a^{*}$, and $a_{t}$ is the arm selected at time step $t$. It is important to note that the agent only receives feedback from the selected arm/action. More specifically, when selecting $a_{t}$ the agent receives a reward $r_{t,a_{t}}$ which is sampled from the underlying reward distribution. As previously mentioned, the agent’s objective is to minimize the cumulative regret, not to learn the complete reward function for each arm. Thus, the agent has to balance delicately between exploration and exploitation.
@@ 2.2 Contextual Bandit Problem
The contextual multi-armed bandit problem is an extension of the classical MAB problem described in Section 2.1. The agent is, at each time step $t$, presented with a context vector $\mathbf{x}_{t,a}\in\mathbb{R}^{d}$ for each action $a\in\mathcal{A}$. For example, a recommendation system could encode a combination of user-related and action-specific features into the context vector $\mathbf{x}_{t,a}$. Then, the expected reward of action $a$ at time $t$ is given by an (unknown) function of the context $q:\mathbb{R}^{d}\rightarrow\mathbb{R}$, such that $\mathbb{E}[r_{t,a}]=q(\mathbf{x}_{t,a})$. Learning to generalize the relationship between the contextual features and the expected reward is crucial for effective learning and minimizing cumulative regret.
@@ 2.3 Combinatorial Bandits
Combinatorial multi-armed bandits (CMAB) (Cesa-Bianchi and Lugosi, 2012) deal with problems where at each time step $t$, a subset $\mathcal{S}_{t}$ (called a super arm) of the set of all base arms $\mathcal{A}$ is selected, instead of an individual arm. The reward of a super arm $\mathcal{S}$ depends on its constituent base arms. When the reward for a super arm is given as a function (e.g., sum) of feedback from its individual base arms (observable if and only if the super arm is selected), it is referred to as semi-bandit feedback. As previously, the evaluation metric is the cumulative regret, and for the combinatorial semi-bandit setting (with the sum of base arm feedback as super arm reward) it can be defined as
@@ 3 Proposed Algorithms
Despite the potential benefits, these types of models have not been studied much for contextual bandits. To the best of our knowledge, there is no previously known work that in a principled way combines tree ensemble models with UCB, which is one of the most effective strategies known for handling the exploration-exploitation dilemma. Further, works that combine tree ensemble methods with Thompson Sampling are limited. We address this research gap and introduce a novel approach for the contextual MAB problem that combines any tree ensemble method with a natural adaption of these two prominent exploration schemes.
Within our framework, we empirically investigate the XGBoost and random forest algorithms and demonstrate promising performance on several benchmark tasks, as well as a combinatorial contextual MAB problem for efficient navigation over a stochastic road network in Section 4. However, we emphasize that our bandit algorithms are sufficiently general to be employed with any decision tree ensemble.
3.1 Tree-Based Weak Learners
@@
In our notation, a tree ensemble regressor $f$ is a collection of $N$ decision trees $\{f_{n}\}$, where the trees are collectively fitted to a set of training samples $\{(\mathbf{x},r)\}$. Each sample consists of a context vector $\mathbf{x}$ and a target value $r$. The fitting procedure can differ between different types of tree ensemble methods, which are often divided into two categories based on the approach used, i.e., bagging or boosting. In both variants, the prediction of each tree is determined by assigning the training samples to distinct leaves, where all samples in the same leaf resemble each other in some way, based on the features attended to in the splitting criteria associated with the ancestor nodes of the leaf. Hence, when fitted to the data, each tree $f_{n}$ receives an input vector $\mathbf{x}$ which it assigns to one of its leaves.
Every leaf is associated with three values that depend on the samples from the training data assigned to that particular leaf. We denote the output of the $n$th tree by $o_{n}$, which is this tree’s contribution to the total ensemble prediction and depends on which leaf the input vector $\mathbf{x}$ is assigned to. The number of training samples assigned to the same leaf as $\mathbf{x}$ in tree $n$ is denoted by $c_{n}$. Finally, the sample variance in the output proposed by the individual training samples assigned to the leaf is denoted $s_{n}^{2}$. This value represents the uncertainty in the output $o_{n}$ associated with the leaf and depends not only on the training samples — but also on the particular tree ensemble method used. We elaborate on the calculations of $s_{n}^{2}$ for two types of tree ensembles in Sections 3.6 and 3.7. Note that, for the sample variance to exist, we must have at least two samples. This is ensured by requiring the tree to be built such that no leaf has fewer than two training samples assigned to it.
The tree ensemble constructs its total target value prediction $p$ by summing up all of the $N$ individual trees’ outputs $o_{n}$:
$p(\mathbf{x})=\sum_{i=1}^{N}o_{i}(\mathbf{x}).$ (3)
The outputs $o_{n}$, in turn, are averages of the $c_{n}$ outputs suggested by each training sample assigned to the same leaf. These suggested outputs are based on the target values of the training samples, with the calculation method depending on the specific type of tree ensemble used. We always consider the full tree ensemble target value predictions as a sum of the individual tree predictions $o_{n}$ to unify the notation of all kinds of tree ensembles. Hence, for averaging models, such as random forests, we assume that the leaf values $o_{n}$ are already weighted by $1/N$.
@@ 3.2 Uncertainty Modeling
Firstly, we assume that the output $o_{n}$ of the $n$’th decision tree, given a context $\mathbf{x}$, is an arithmetic average of $c_{n}$ independent and identically distributed random variables with finite mean $\mu_{n}$ and variance $\sigma_{n}^{2}$. By this assumption, $o_{n}$ is itself a random variable with mean $\mu_{n}$ and variance $\frac{\sigma_{n}^{2}}{c_{n}}$. Also, we can approximate $\mu_{n}$ and $\sigma_{n}^{2}$ by the sample mean and variance respectively, based on the training samples assigned to the leaf. Moreover, the central limit theorem (CLT) (see e.g., Dodge, 2008) ensures that, as $c_{n}\to\infty$, the distribution of $o_{n}$ tends to a Gaussian distribution. See Appendix A.2 for more details about the independence assumption.
Secondly, if we also assume the output of each one of the $N$ trees in the ensemble to be independent of every other tree, we have that the total tree ensemble’s target value prediction, which is a sum of $N$ random variables, is normally distributed with mean $\mu=\sum_{i=1}^{N}\mu_{i}$ and variance $\sigma^{2}=\sum_{i=1}^{N}\sigma_{i}^{2}$.
@@ 3.3 Tree Ensemble Upper Confidence Bound
UCB methods act under the principle of optimism in the face of uncertainty, and have been established as some of the most prominent approaches to handling exploration in bandit problems. A classic example is the UCB1 algorithm (Auer et al., 2002) which builds confidence bounds around the expected reward from each arm depending on the fraction of times it has been played, and for which there are proven upper bounds on the expected regret. One disadvantage of the method, however, is that it does not take the variance of the observed rewards into account, which may lead to a sub-optimal exploration strategy in practice. In light of this, Auer et al. (2002) further proposed the UCB1-Tuned and UCB1-Normal algorithms, which extend UCB1 by including variance estimates, and demonstrate better performance experimentally. The main difference between the two extended versions is that UCB1-Tuned assumes Bernoulli-distributed rewards, while UCB1-Normal is constructed for Gaussian rewards. At each time step $t$, UCB1-Normal selects the arm $a$ with the maximal upper confidence bound $U_{t,a}$, calculated as
$U_{t,a}\leftarrow\tilde{\mu}_{t,a}+\sqrt{16\tilde{\sigma}^{2}_{t,a}\frac{\ln(t-1)}{m_{t,a}}},$ (4)
where $m_{t,a}$ is the number of times arm $a$ has been played so far, and $\tilde{\mu}_{t,a}$ and $\tilde{\sigma}^{2}_{t,a}$ are the sample mean and variance of the corresponding observed rewards, respectively. For the sample variance to be defined for all arms, they must have been played at least twice first.
By the assumptions we have made on the tree ensembles, the predictions of the expected rewards will be approximately Gaussian. Therefore, we propose an algorithm called Tree Ensemble Upper Confidence Bound (TEUCB) in Algorithm 1, which draws inspiration from UCB1-Normal, and is suitable for both Gaussian and Bernoulli bandits. As seen in lines 27 and 31 of Algorithm 1, the selection rule of TEUCB closely resembles that of UCB1-Normal, but is constructed specifically for contextual MABs where the arms available in each time step are characterized by their context vectors. Therefore, TEUCB considers each individual contribution from all samples in a leaf as a sample of that leaf’s output distribution, which is demonstrated in line 14. The total prediction for a context $\mathbf{x}_{t,a}$ is subsequently computed as the sum of sampled contributions from each leaf that $\mathbf{x}_{t,a}$ is assigned to in the ensemble (lines 16, 17, 18).
Algorithm 1 Tree Ensemble Upper Confidence Bound / Tree Ensemble Thompson Sampling
1:  Input: Number of rounds $T$, number of initial random selection rounds $T_{I}$, number of trees in ensemble $N$, exploration factor $\nu$, tree ensemble regressor $f$, bandit method (UCB or TS).
@@
7:     Fit tree ensemble $f$ to previously observed context-reward pairs $\left\{(\mathbf{x}_{i,a_{i}},r_{i,a_{i}})\right\}_{i=1}^{t-1}$ and set leaf values
8:     Observe current contexts $\{\mathbf{x}_{t,a}\}_{a=1}^{K}$
9:     for $a=1$ to $K$ do
10:        Initialize arm parameters:$\tilde{\mu}_{t,a}\leftarrow 0,\quad\tilde{\sigma}^{2}_{t,a}\leftarrow 0,\quad c_{t,a}\leftarrow 0$
11:        for n = 1 to $N$ do
12:           Assign leaf values:
13:           $(o_{t,a,n},s_{t,a,n},c_{t,a,n})\leftarrow f_{n}\left(\mathbf{x}_{t,a}\right)$
14:           Update leaf parameters:$\tilde{\mu}_{t,a,n}\leftarrow o_{t,a,n},\quad\tilde{\sigma}^{2}_{t,a,n}\leftarrow\frac{s^{2}_{t,a,n}}{c_{t,a,n}}$
15:           Increment arm parameters:
16:           $\tilde{\mu}_{t,a}\leftarrow\tilde{\mu}_{t,a}+\tilde{\mu}_{t,a,n}$
17:           $\tilde{\sigma}^{2}_{t,a}\leftarrow\tilde{\sigma}^{2}_{t,a}+\tilde{\sigma}^{2}_{t,a,n}$
18:           $c_{t,a}\leftarrow c_{t,a}+c_{t,a,n}$
19:        end for
20:        if bandit method is UCB then
21:           $U_{t,a}\leftarrow\tilde{\mu}_{t,a}+\sqrt{\nu^{2}\tilde{\sigma}^{2}_{t,a}\frac{\ln(t-1)}{c_{t,a}}}$
22:        else if bandit method is TS then
23:           $\tilde{r}_{t,a}\sim\mathcal{N}(\tilde{\mu}_{t,a},\nu^{2}\tilde{\sigma}^{2}_{t,a})$
24:        end if
25:     end for
26:     if bandit method is UCB then
27:        $a_{t}\leftarrow\text{argmax}_{a}U_{t,a}$
28:     else if bandit method is TS then
29:        $a_{t}\leftarrow\text{argmax}_{a}\tilde{r}_{t,a}$
30:     end if
31:     Play $a_{t}$ and observe reward $r_{t,a_{t}}$
32:  end for